OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 100. Find y when x = 6

OpenStudy (anonymous):

y=kx^2

OpenStudy (anonymous):

kx^2 = y k5^2 = 100 k = 4 4(6^2) =y y =144

OpenStudy (anonymous):

k=4, thus y=36*4=144

OpenStudy (anonymous):

where does the 4 come from though?

OpenStudy (anonymous):

show the work please so I can understand it.

OpenStudy (jamesj):

If k.x^2 = y and x = 5, y = 100, then k.5^2 = 100 i.e., 25k = 100 i.e., k = 4